If two objects are at a very large distance from each other ad move towards each other due to mutual forces then what is the indiviual kinetic energy considering each to be of mass m and final distance r.


plz answer

Grade:

1 Answers

vikas askiitian expert
13 years ago

accordin to conservation of linear momentam

mv1 + mv2 = final momentam = 0 = initial momentam

v1 = -v2

(KE)1 = (KE)2           (at any time)

 

now using conservation of energy

PE + KE = initial = 0 + 0 = 0

PE + KE = final = -Gm2/r + (KE)1 + (KE)2

(KE)1 = (KE)1 = KE so

KE = Gm2/2r         ................1

kinetic energy of both will be Gm2/2r